OpenStudy (mousejkl):

Carrie pays a monthly fee of $8 to stream movies online.  She also pays a fee every time she rents a movie. Last month Carrie rented 6 movies, and her total bill for the month was $32. Select from the drop-down menu to correctly complete each statement. The monthly fee for Carrie to stream movies is . Carrie also pays every time she rents a movie.

OpenStudy (cp123):

8+x=32. The monthly fee is $8, it is fixed so there is no variable since you are solving for this month only. solve for x to see how much each rental costs.

OpenStudy (cp123):

oops i meant 6x, since she rented 6 movies for an undefined price.

OpenStudy (cp123):

8+6x=32. Now solve for x. so you need to isolate x on its own. start by subtracting 8 from each side. 6x=24. Now, since 6 is being multiplied by x, you need to divide both sides by 6.

OpenStudy (cp123):

So, since 24/6=4, x=4. She pays $4 for each movie she rents.

OpenStudy (cp123):

To check the answer, replace 4 in the original equation for x. 8+6(4)=32. 8+24=32. 32=32! this is a true statement, so the answer is correct